OpenStudy (anonymous):

Differentiate: y = (v^3 - 2v[sqrt[v]])/v

OpenStudy (bahrom7893):

split this up into two fractions, it will become a piece of cake

OpenStudy (bahrom7893):

y = v^2 - 2sqrt(v) y' = 2v - 2*(1/2)*v^(-1/2) = 2x - 1/sqrt(v)

OpenStudy (anonymous):

yeah i got (2v^2 - v^(3/2))/v^2.... so from there how does that equate to 2v - 1/sqrt(v) ?

OpenStudy (bahrom7893):